What are two assumptions to analyze the motion, if the ball moves in a curved path? A particle starts from the origin at t=0 with an initial velocity having an x component of 20 m/s and a y component of – 15 m/s.The particle moves in the xy plane with an x component f acceleration only, gives by a_x= 4.0 m/s^2.Calculate the velocity and speed of particle at t = 5 s. *


1
Expert's answer
2021-05-19T11:03:32-0400

By definition, the velocity of the particle under a constant acceleration is:


"\mathbf{v} = \mathbf{v}_0 + \mathbf{a}t"

where "\mathbf{v}_0 = (20, -15)\space m/s" is the vector of the initial velocity, "\mathbf{a} = (4,0)\space m/s^2" is the vector of the acceleration, and "t =5s" is the time. Thus, obtain:


"\mathbf{v} = (20, -15)\space m/s \space +\space (4,0)\space m/s^2\cdot 5s = (40,-15)\space m/s"

The speed is the magnitude of the velocity:


"v = \sqrt{(40)^2 + (-15)^2} \approx 42.7m/s"

Answer. The velocity is "(40,-15)\space m/s"., the speed is 42.7 m/s.